find the missing coordinate give the slope of the line (2,3) and (x,9) m= 3/2
thanks

Answer by lwsshak3(11628)

find the missing coordinate give the slope of the line (2,3) and (x,9) m= 3/2
Slope,m=∆y/∆x=(9-3)/x-2=3/2
6/(x-2)=3/2
12=3x-6
3x=18
x=6